(a) Monty Co. sold $2,030,000 of 10%, 10-year bonds at 105 on January 1, 2017. The bonds were dated January 1, 2017, and pay interest on July 1 and January 1. If Monty uses the straight-line method to amortize bond premium or discount, determine the amount of interest expense to be reported on July 1, 2017, and December 31, 2017. (Round answer to 0 decimal places, e.g. 38,548.)


1) What is the interest expense to be recorded?

(b) Flounder Inc. issued $660,000 of 8%, 10-year bonds on June 30, 2017, for $577,750. This price provided a yield of 10% on the bonds. Interest is payable semiannually on December 31 and June 30. If Flounder uses the effective-interest method, determine the amount of interest expense to record if financial statements are issued on October 31, 2017. (Round intermediate calculations to 6 decimal places, e.g. 1.251247 and final answer to 0 decimal places, e.g. 38,548.)

2) What is the interest expense to be recorded?
Answer #1
a) Bonds issue price (2,030,000*1.05)= 2131500
Bonds face value 2,030,000
Bonds premium 101,500
Amortization of premium = 101,500/20
5075
interest expense to be reported on july and December
interest paid (2,030,000*5%)= 101500
less:premium on bonds payable amortized 5,075
interest expense to be reported on july and December 96,425 answer
b) interest expense to be recorded on Oct 31,2017
interest expense = 577,750*10%*4/12
19258 answer